### v1

# Understanding uncertainty in Bayesian cluster analysis

Abstract
The Bayesian approach to clustering is often appreciated for its ability to provide uncertainty in the partition structure. However, summarizing the posterior distribution over the clustering structure can be challenging, due the discrete, unordered nature and massive dimension of the space. While recent advancements provide a single clustering estimate to represent the posterior, this ignores uncertainty and may even be unrepresentative in instances where the posterior is multimodal. To enhance our understanding of uncertainty, we propose a WASserstein Approximation for Bayesian clusterIng (WASABI), which summarizes the posterior samples with not one, but multiple clustering estimates, each corresponding to a different part of the space of partitions that receives substantial posterior mass. Specifically, we find such clustering estimates by approximating the posterior distribution in a Wasserstein distance sense, equipped with a suitable metric on the partition space. An interesting byproduct is that a locally optimal solution to this problem can be found using a k-medoids-like algorithm on the partition space to divide the posterior samples into different groups, each represented by one of the clustering estimates. Using both synthetic and real datasets, we show that our proposal helps to improve the understanding of uncertainty, particularly when the data clusters are not well separated or when the employed model is misspecified.
1 Introduction
Clustering is one of the canonical forms of unsupervised learning and has numerous applications in various contexts, from finance and economics (Saunders, 1980), to natural language processing and computer science (Blei et al., 2003), biomedicine (Oyelade et al., 2016; Chauvel et al., 2019) and more. It aims to discover groups of similar data points and characterize the patterns within. While algorithmic schemes, such as k-means (Hartigan and Wong, 1979; Jain, 2010), are widely used, only a single clustering solution is provided, without any measure of uncertainty in the solution. To overcome this, the Bayesian approach is a natural choice. Through Bayes rule, the prior over the space of clusterings is combined with the model (or loss) to assess homogeneity within group, producing posterior beliefs and uncertainty for all possible clustering structures and patterns within each cluster.
However, understanding this posterior over the clustering structure can be daunting, due to the huge dimension and discrete, unordered nature of the space of partitions. Indeed, the number of clusters is rarely known in practice, and many Bayesian methods account for this uncertainty, further increasing the dimension of the space. For example, in a Bayesian nonparametric setting, allowing anywhere between 1 and $n$ clusters, the total number of ways to partition the data is a Bell number, which is $B_{20}=51,724,158,235,372$ for only 20 data points. Even if the number of clusters is restricted in a parametric setting, the number of possible partitions is still massive, e.g. restricting to at most 5 clusters, there are still $795,019,337,135$ possible clusterings of $20$ data points.
Clearly, this makes any exact posterior computations infeasible, and to overcome this, Markov chain Monte Carlo (MCMC) is the most common tool. MCMC provides tens of thousands of clustering solutions, representing (asymptotically exact) draws from the posterior. Due to high dimension of the space, almost all of these clustering solutions are unique and describing the behavior and uncertainty across them is far from straightforward. Alternative inference procedures can also be used, such as importance sampling (Zhang and Williamson, 2019), sequential Monte Carlo (MacEachern et al., 1999; Härkönen et al., 2022), or posterior bootstrap (Fong et al., 2019), which similarly provide a very large number of different clustering solutions. An increasingly used alternative to posterior simulation is approximate optimization-based techniques, such as variational inference. However, variational inference is well known to suffer from mode collapse, with the approximate posterior concentrating around a single clustering solution. Thus, it tends to provide unsatisfying and unrealistic measures of certainty in the clustering structure. A simple approach to overcome this combines multiple variational approximations through ensembles or stacking (Yao et al., 2022), which similarly produces a large number of different clustering solutions. Ensembles are also used to combat poor mixing of MCMC in high-dimensions through consensus clustering (Coleman et al., 2022).
When faced with this challenge of describing the numerous clustering solutions produced by these inference algorithms, a first task is to find a single representative clustering. From a decision-theoretic viewpoint, the optimal Bayes estimator of the clustering is found by minimizing the posterior expected loss. This requires the choice of an appropriate loss on the space of partitions, and the 0-1 loss is a simple choice, resulting in the maximum a posteriori (MAP) partition estimator. However, the 0-1 loss between two partitions is zero if and only if the two partitions coincide, and it equally penalizes any two different partitions, irrespective of how similar or different they are. This property makes the 0-1 loss quite unappealing. Thus, more general losses are required and have been used, such as Binder’s loss (Binder, 1978; Dahl, 2006; Lau and Green, 2007), the adjusted Rand Index (ARI, Fritsch and Ickstadt, 2009), the variation of information (VI, Meilă, 2007; Wade and Ghahramani, 2018), the normalized VI (Vinh et al., 2010; Rastelli and Friel, 2018), the generalized Binder’s loss and VI (Dahl et al., 2022), among others (Quintana and Iglesias, 2003; Buch et al., 2024; Nguyen and Mueller, 2024). However, these methods all focus on finding a single optimal clustering estimate (a critical aspect that is also highlighted in the discussion of (Rastelli and Friel, 2018; Dahl et al., 2022)).
Beyond providing a single clustering solution, a main motivation for the Bayesian approach is the additional uncertainty in the clustering structure. Thus, tools are also required to visualize and describe this uncertainty, which again is difficult due to the sheer number of partitions, combined with the variable number of clusters and label-switching. One important tool is the posterior similarity matrix, whose elements represent the posterior probability that a pair of data points are clustered together (which is also referred to as the co-clustering or consensus matrix). Additionally, Wade and Ghahramani (2018) developed credible balls to characterize uncertainty around the clustering estimate, but accurate estimation of the bounds of the credible ball based on the MCMC draws can be challenging in certain cases due to massive dimension of the partition space. An alternative greedy approach to compute the credible bounds, that allows searching beyond the MCMC draws, is presented in Buch et al. (2024). Another approach (Lavigne and Liverani, 2024) approximates and visualizes uncertainty in the allocation of each data point, but only provides a limited view of uncertainty by conditioning on the allocation of all other points. Lastly, Do et al. (2024) represent uncertainty in the underlying mixing measure for Gaussian mixtures through a dendogram.
(a) Posterior on partitions
(b) Posterior similarity matrix
(c) Histogram of the data
(d) WASABI summarizes with multiple, weighted clustering estimates
Figure 1: Slightly bimodal example. The MAP partition of a DPM for data generated from a slightly bimodal mixture (Figure 1(c)) contains only a single cluster, yet the posterior similarity matrix (Figure 1(b)) suggests uncertainty in additional clusters. To further examine the posterior on the space of partitions, Figure 1(a) shows the log posterior relative to the MAP against the VI distance to the MAP, where each point represents a partition colored by its number of clusters, with cluster sizes reported for some. Multimodality in the posterior on partitions is evident with one mode corresponding to the MAP and another around partitions with two clusters of more equal size. WASABI summarizes with multiple, weighted partitions (Figure 1(d)), reflecting the two different modes of clustering in this example.
In this work, we contribute an important novel tool to enhance our understanding of uncertainty in the posterior on partitions, that is applicable to the wide range of Bayesian clustering models for diverse data types. As an illustrative toy example, consider data generated from a slightly bimodal mixture of Gaussians (Figure 1(c)). Interestingly, when focusing on the marginal posterior of the number of clusters, a number of papers Miller and Harrison (2013, 2014); Ohn and Lin (2020); Ascolani et al. (2022); Alamichel et al. (2024) have demonstrated inconsistency and overestimation of the number of clusters in Bayesian nonparametric mixture models, such as the popular Dirichlet process mixture (DPM), if the hyperparameters are not appropriately tuned. Despite this, the optimal clustering estimate can still recover the true clustering structure under an appropriate choice of loss (see (Rajkowski, 2019) for a theoretical analysis of the MAP and Wade (2023) for an empirical analysis comparing the MAP, VI and Binder’s loss). However, when the clusters are not well separated, the different estimators often disagree; indeed the MAP and VI clustering estimate only contain a single cluster [^1] While previous work found agreement between the clustering solutions which minimize the expected VI and its more computationally efficient lower bound, in this slightly bimodal example, they disagree, with the latter reporting two clusters., while the ARI and Binder’s contain two clusters in the slightly bimodal example (Figure 1(c)). The posterior similarity matrix (Figure 1(b)) also suggests possible additional clusters, but without a clear separation between groups. Further examination of the posterior on partitions (Figure 1(a)) shows multimodality in the posterior. Starting from the MAP partition with a single cluster, the posterior mass decreases slightly as a single data point is split off, and continues to decrease as we move further from the MAP, creating additional small clusters. But, it starts to increase again around more distant partitions with two clusters of roughly equal size.
To capture this multimodality and describe the main clustering patterns present across the posterior, we propose to summarize the posterior with not one, but multiple optimal partitions, found by minimizing the Wasserstein distance equipped with an appropriate loss. Our WASserstein Approximation for Bayesian clusterIng (WASABI) also provides weights attached to the optimal partitions (referred to as particles) and reflects the different modes of clustering in the slightly bimodal example (Figure 1(d)). We focus on the VI loss, due to its benefits shown in (e.g Wade and Ghahramani, 2018; Rastelli and Friel, 2018; Dahl et al., 2022), but other losses can also be considered. This builds on the idea of Balocchi et al. (2023), who proposed an ensemble optimization method that approximates the posterior distribution by targeting several partition estimates through minimization of the Kullback-Leibler divergence. Instead, we use the Wasserstein distance, motivated by its role in hierarchical modeling and clustering. Indeed, it has been shown to be the natural distance metric for convergence theory of mixing measures (Nguyen, 2013, 2016; Ho and Nguyen, 2016; Guha et al., 2021). Furthermore, the Wasserstein distance is one of the few distances that allows measuring the discrepancy between distributions based on an arbitrary metric on the space on which they are supported. An interesting byproduct of using the Wasserstein distance is that the problem of determining the optimal discrete distribution with minimal Wasserstein distance from the empirical distribution can be directly linked to k-mediods clustering (Pollard, 1982; Graf and Luschgy, 2000; Canas and Rosasco, 2012). As a consequence, the tool we propose can be seen as employing k-mediods clustering to divide the samples into distinct groups and using the mediods as optimal partitions summarizing the multiple modes of the posterior. This will greatly enhance researchers’ toolbox to better describe and communicate findings and showcase the benefits of the uncertainty provided by the Bayesian approach to clustering. The code implementing the WASABI algorithm is available as an R package at https://github.com/cecilia-balocchi/WASABI.
2 Review
In the context of clustering, the observed data consists of measurements $\bm{y}=(y_{1},\ldots,y_{n})$ drawn from a heterogeneous population consisting of an unknown number homogeneous sub-populations. The observed $y_{i}\in\mathcal{Y}$ may be continuous, discrete, mixed, or more complex in nature (e.g. functional data). Each data point is associated with a discrete latent variable $c_{i}$ (also called the allocation variable) indicating the group membership of the data point, i.e. $c_{i}=j$ if $y_{i}$ belongs to the $j$th group, and $c_{i}=c_{i^{\prime}}$ if $y_{i}$ and $y_{i^{\prime}}$ belong to the same sub-population. We are interested in obtaining estimates and understanding the uncertainty of clustering structure characterized by the latent $\rho_{n}=(c_{1},\ldots,c_{n})$. To achieve this, the Bayesian approach constructs a posterior distribution over the space of partitions,
$\pi(\rho_{n}|\bm{y})\propto p(\bm{y}\mid\rho_{n})\pi(\rho_{n}),$ (1)
where $\pi(\rho_{n})$ represents the prior over the space of partitions and $p(\bm{y}\mid\rho_{n})$ is the likelihood in a model-based approach or defined based on a loss function in a loss-based approach (Rigon et al., 2020). For reviews on Bayesian clustering, we refer the reader to Wade (2023); Grazian (2023).
It is worth emphasizing that clustering is often referred to as an ill-posed problem, as it aims to discover unknown patterns or structures in the data. The notion of a cluster depends on the application at hand and can often be challenging to characterize formally. A unique clustering solution often does not exist (Hennig, 2015). Thus, one must carefully consider the model or loss employed and importantly, also characterize uncertainty in the clustering solution. To achieve the latter, Bayesian cluster analysis provides a formal framework through both the posterior distribution over the entire space of clusterings and by creating an ensemble of clustering solutions sampled from the posterior. Moreover, this also helps to mitigate sensitivity to local optimum which adversely impact all clustering algorithms due to the sheer size of the space.
2.1 Random Partition Models
A key ingredient in the Bayesian approach is the prior $\pi(\rho_{n})$ over the clustering structure. This is referred to as random partition model, as it involves randomly assigning observations to clusters, or equivalently, randomly partitioning the indices of the data points $\{1,\ldots,n\}$ into $K_{n}$ non-empty and mutually exhaustive sets $C_{j}$ for $j=1,\ldots,K_{n}$, where $K_{n}$ represents the number of clusters. Thus, the partition $\rho_{n}$ can also be represented as $\rho_{n}=\{C_{1},\ldots,C_{K_{n}}\}$, where the sets satisfy $C_{j}\cap C_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $C_{1}\cup\ldots\cup C_{K_{n}}=\{1,\ldots,n\}$; essentially, each $C_{j}$ contains the indices of data points in the $j$th cluster. For ease of notation, the sample size is dropped from $\rho_{n}$ and $K_{n}$, when the context is clear.
In practice, the number of clusters $K$ is rarely known, and choosing this number is an important and difficult concern in cluster analysis. While information criteria or model selection tools, such as the Bayesian information criterion (BIC), are commonly used, this ignores uncertainty in the number of clusters and disregards information from the numerous models fits using different choices of $K$. Instead, our focus is approaches that account for this uncertainty, namely, through 1) mixtures of finite mixtures, which naturally incorporate a prior on the unknown $K$ Miller and Harrison (2018); Nobile and Fearnside (2007); Richardson and Green (1997); 2) sparse overfitted mixtures that specify an upper bound on the number of clusters and encourage extra components to be emptied through sparsity promoting priors (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016; Van Havre et al., 2015; Frühwirth-Schnatter et al., 2021); and 3) Bayesian nonparametric (BNP) mixtures (Müller, 2019) that allow the number of clusters to grow unboundedly with the data, with the Dirichlet process (DP) mixture (Lo, 1984) being the most-widely used example. For example, letting $L$ denote an upper bound on the number of clusters and $\gamma>0$, the random partition model for the overfitted mixture (corresponding to a Dirichlet prior on the weights of the mixture with $L$ components) is
$\pi(\rho)$ $=\frac{\Gamma(\gamma L)L!}{\Gamma(\gamma L+n)(L-K)!}\prod^{K}_{j=1}\frac{\Gamma(n_{j}+\gamma)}{\Gamma(\gamma)},$
with $n_{j}=|C_{j}|$ denoting the cluster size. In general, random partition models specify priors on the discrete, partially-ordered space of partitions, denoted by $\mathcal{P}$, which has massive dimension, even more so when accounting for the unknown number of clusters.
2.2 Inference
The massive dimension makes computing the posterior in eq. (1) infeasible. Instead, MCMC is commonly used to obtain asymptotically exact draws, denoted by $\rho^{(t)}$ for $t=1,\ldots,T$. The number of draws is typically in the tens to hundreds of thousands, and the $\rho^{(t)}$ are mostly unique (due to the size of the space), vary in the number of clusters, and suffer from label-switching.
To provide a single, representative clustering solution, the optimal Bayes estimator minimizes the posterior expected loss. For a partition-specific loss, Wade and Ghahramani (2018) recommend the Variation of Information (VI) Meilă (2007), which defines the distance between any two partitions $\rho_{1}$ and $\rho_{2}$ as:
$\text{VI}(\rho_{1},\rho_{2})$ $=\text{H}(\rho_{1})+\text{H}(\rho_{2})-2\text{MI}(\rho_{1},\rho_{2})$
$=-\sum_{j_{1}=1}^{K_{1}}\frac{n_{j_{1},+}}{n}\log_{2}\left(\frac{n_{j_{1},+}}{n}\right)-\sum_{j_{2}=1}^{K_{2}}\frac{n_{+,j_{2}}}{n}\log_{2}\left(\frac{n_{+,j_{2}}}{n}\right)-2\sum_{j_{1}=1}^{K_{1}}\sum_{j_{2}=1}^{K_{2}}\frac{n_{j_{1},j_{2}}}{n}\log_{2}\left(\frac{n_{j_{1},j_{2}}n}{n_{j_{1},+}n_{+,j_{2}}}\right),$
where the first two terms represent the entropy of the two partitions and the last term is the mutual information between the partitions. The counts $n_{j_{1},j_{2}}$ represent the cross-tabulation between the partitions, i.e. $n_{j_{1},j_{2}}$ is the number of data points in cluster $j_{1}$ in $\rho_{1}$ and cluster $j_{2}$ in $\rho_{2}$, and $K_{1}$ and $K_{2}$ are the number of clusters in $\rho_{1}$ and $\rho_{2}$, respectively. Thus, the Bayes estimator under the VI loss (called the minVI partition) is:
$\rho^{*}=\min_{\rho\in\mathcal{P}}\mathbb{E}\left[\text{VI}(\rho,\rho^{\prime})\mid\bm{y}\right]\approx\min_{\rho\in\mathcal{P}}\frac{1}{T}\sum_{t=1}^{T}\text{VI}(\rho,\rho^{(t)}).$
While direct minimization cannot be performed, effective algorithms exist Wade and Ghahramani (2018); Rastelli and Friel (2018); Dahl et al. (2022).
3 Methods
Summarizing the posterior with a single optimal partition not only disregards uncertainty in the clustering structure, but can also be misleading in the case of multimodality (see Example 2). To address these issues, we propose to summarize with not one, but multiple optimal partitions. Our approach borrows ideas from Balocchi et al. (2023) to approximate the posterior $\pi(\rho|\bm{y})$ with a simpler discrete distribution $q=\sum_{\ell=1}^{L}w_{\ell}\delta_{\rho_{\ell}}$ supported on a small number $L$ of partitions. Balocchi et al. focus on minimizing the Kullback-Leibler (KL) divergence and show that the optimal distribution $q$ is supported on the set of $L$ partitions that have largest posterior mass (for $L=1$, this reduces to the MAP partition). However, such partitions tend to be very similar to each other, with only one or two points having a different cluster allocation. For example, in Figure 1(a), even with the choice of $L=501$, the support of $q$ consists of the partition with a single large cluster of $500$ points and all partitions with one large cluster of size $499$ and one singleton. This would only summarize uncertainty around the mode, not the variability in the posterior.
Instead, we replace the KL divergence with the Wasserstein distance, which allows the use of metrics on the space of partitions. Here, we focus on the VI metric, but this can be extended to other choices. To formally define the Wasserstein-VI distance, let $(\mathcal{P},d_{\text{VI}})$ be the metric space on the set of partitions $\mathcal{P}$ embedded with the VI metric $d_{\text{VI}}$. Let $p$ and $q$ represent two distributions on the space of partitions $(\mathcal{P},d_{\text{VI}})$. If $\mathcal{J}(p,q)$ is the set of couplings of $p$ and $q$, i.e. the collection of distributions $J(\rho,\rho^{\prime})$ on $\mathcal{P}\times\mathcal{P}$ with marginals $p$ and $q$ on the first and second factor respectively, then we define the Wasserstein distance as
$W_{\text{VI}}(p,q)=\inf_{J\in\mathcal{J}(p,q)}\sum_{\rho,\rho^{\prime}\in\mathcal{P}}d_{\text{VI}}(\rho,\rho^{\prime})J(\rho,\rho^{\prime}).$
Denote with $\mathcal{Q}_{L}=\{\sum_{\ell=1}^{L}w_{\ell}\delta_{\rho_{\ell}}:\sum_{\ell=1}^{L}w_{\ell}=1,w_{\ell}\geq 0,\rho_{\ell}\in\mathcal{P},\textrm{ for }\ell=1,\ldots,L\}$ the collection of discrete distributions supported on $L$ points.
Definition 1.
The WASABI posterior is the discrete distribution $q^{*}=\sum_{\ell=1}^{L}w^{*}_{\ell}\delta_{\rho^{*}_{\ell}}\in\mathcal{Q}_{L}$ that best approximates the posterior in a Wasserstein-VI sense,
$q^{*}=\operatornamewithlimits{argmin}_{q\in\mathcal{Q}_{L}}W_{\text{VI}}(\pi(\cdot|\bm{y}),q(\cdot)).$ (2)
The quality of the approximation is quantified by the Wasserstein-VI distance, $W_{\text{VI}}(\pi(\cdot|\bm{y}),q^{*}(\cdot))$, which measures minimal amount of uncertainty lost when summarizing $L$ partitions.
Remark.
In the special case of $L=1$, the Wasserstein-VI distance reduces to the posterior expected VI. Thus, the optimal distribution $q^{*}$ is a point mass supported on the partition $\rho^{*}$ that minimizes the posterior expected VI (the minVI partition). The quality of the approximation and spread around $\rho^{*}$ is quantified by the posterior expected VI of $\rho^{*}$.
In the general case of $L>1$, we can rewrite the problem of finding the WASABI posterior using properties of the Wasserstein distance between two discrete distributions.
Proposition 1.
The WASABI posterior $q^{*}$, which is the solution to (2), is found by identifying a set of “centers” or “particles” $\bm{\rho}^{*}=\{\rho^{*}_{1},\ldots,\rho^{*}_{L}\}$ that minimize
$\sum_{\ell=1}^{L}\sum_{\rho\in\mathcal{N}_{\ell}}d_{\text{VI}}(\rho,\rho^{*}_{\ell})\pi(\rho|\bm{y}),$ (3)
where $\mathcal{N}_{\ell}=\{\rho:d_{\text{VI}}(\rho,\rho^{*}_{\ell})<d_{\text{VI}}(\rho,\rho^{*}_{\ell^{\prime}})\textrm{ for all }\ell^{\prime}\neq\ell\}$ corresponds to the set of partitions that are closer to $\rho^{*}_{\ell}$ than to any other center. We shall name $\mathcal{N}_{\ell}$ the “region of attraction” of the center $\rho^{*}_{\ell}$. The optimal weights $w^{*}_{\ell}$ associated to each center are found by $w^{*}_{\ell}=\sum_{\rho\in\mathcal{N}_{\ell}}\pi(\rho|\bm{y}).$
In other words, the WASABI posterior is supported on the centers $\rho^{*}_{\ell}$ that minimize (3), and the probability $w^{*}_{\ell}$ associated with each of them is equal to the posterior mass of its region of attraction. See Appendix A for a proof of Proposition 1. Note that this result is known as a general property of the Wasserstein distance, see for example Section 3.1 of Cuturi and Doucet (2014).
In practice, even for moderate $n$, the objective function in (3) is intractable due to the dimension of $\mathcal{P}$. A similar approximation of the posterior can however be considered when an estimate of the posterior distribution is provided through MCMC samples. Let $\hat{\pi}(\rho|\bm{y})$ be the approximation to the posterior provided by the MCMC draws, $\hat{\pi}(\rho|\bm{y})=\frac{1}{T}\sum_{t=1}^{T}\delta_{\rho^{(t)}}$. We can then approximate $\hat{\pi}(\rho|\bm{y})$ with $q$ in a Wasserstein-VI sense.
Definition 2.
Given a set $\{\rho^{(1)},\ldots,\rho^{(T)}\}$ of MCMC draws from the posterior $\pi(\cdot|\bm{y})$, the WASABI posterior is the discrete distribution $q^{*}=\sum_{\ell=1}^{L}w^{*}_{\ell}\delta_{\rho^{*}_{\ell}}\in\mathcal{Q}_{L}$ that is closest in a Wasserstein-VI sense to the empirical approximation of the posterior, $\hat{\pi}(\rho|\bm{y})=\frac{1}{T}\sum_{t=1}^{T}\delta_{\rho^{(t)}}$,
$q^{*}=\operatornamewithlimits{argmin}_{q\in\mathcal{Q}_{L}}W_{\text{VI}}(\hat{\pi}(\cdot|\bm{y}),q(\cdot)).$ (4)
Again, thanks to the properties of the Wasserstein distance between two discrete distributions, we can find the WASABI posterior by reframing the problem as follows.
Proposition 2.
The WASABI posterior $q^{*}\in\mathcal{Q}_{L}$, which is the solution to (4), is supported on the “centers” or “particles” $\bm{\rho}^{*}=\{\rho^{*}_{1},\ldots,\rho^{*}_{L}\}$ which minimize
$\frac{1}{T}\sum_{\ell=1}^{L}\sum_{\rho^{(t)}\in\widehat{\mathcal{N}}_{\ell}}d_{\text{VI}}(\rho^{(t)},\rho^{*}_{\ell}),$ (5)
where $\widehat{\mathcal{N}}_{\ell}=\{\rho^{(t)}:d_{\text{VI}}(\rho^{(t)},\rho^{*}_{\ell})<d_{\text{VI}}(\rho^{(t)},\rho^{*}_{\ell^{\prime}})\textrm{ for all }\ell^{\prime}\neq\ell\}$; we refer to $\widehat{\mathcal{N}}_{\ell}$ as the “empirical region of attraction” of $\rho^{*}_{\ell}$. The optimal weights $w^{*}_{\ell}$ associated to each center are proportional to the size of $\widehat{\mathcal{N}}_{\ell}$, i.e. $w^{*}_{\ell}=|\widehat{\mathcal{N}}_{\ell}|$ / T.
As described by Canas and Rosasco (2012) and first suggested by Pollard (1982), an interesting connection arises between the k-means algorithm and the approximation in a Wasserstein sense (equipped with the Euclidean distance) of an empirical distribution with a discrete distribution supported on $k$ points. In particular, the k-means algorithm provides a locally optimal solution to this problem. In our framework, we can obtain a locally optimal set of partitions that minimize (5) by applying a k-medoids-like algorithm, which generalizes k-means beyond the Euclidean distance, to the set of MCMC samples $\{\rho^{(t)}:t=1,\ldots,T\}$, with $k=L$ and distance equal to the VI distance. In other words, the MCMC-sampled partitions are clustered into $L$ groups, and each of these groups is represented by one of the centers $\rho^{*}_{\ell}$. These same centers are the support of the WASABI posterior $q^{*}$, with weights determined by the size of each group. While clustering a set of partitions may seem tautological, the value of this approach lies in its ability to summarize a large set of discrete objects with a smaller (but greater than one) number of points, each of which represents a different group or “cluster” of partitions.
Example 1.
Consider the slightly bimodal example from the Introduction (Figure 1), with data generated from an equal-weight mixture of two unit-variance Gaussians (means $+1.1$ and $-1.1$). The minVI estimator identifies only one cluster, yet the posterior similarity matrix (PSM) suggests a possible two-cluster structure. Approximating the posterior with WASABI (using two particles), we find one particle corresponding to the single-cluster partition ($w_{1}\approx 0.7$), and another recovers the two similarly-sized clusters. In this sense, WASABI is able to recover the bimodal structure of the posterior, visible in Figure 1(a).
3.1 Algorithm
Algorithm 1 Find the approximate WASABI posterior
Input: MCMC samples $\{\rho^{(t)}\}_{t=1}^{T}$, number of particles $L$, initialization method init, tolerance $\epsilon$
Initialize $\rho^{*}_{1:L}$ using method init
repeat
     N-update step:
     Compute $\text{VI}(\rho^{(t)},\rho^{*}_{\ell})$ for all $t$ and $\ell$
     Assign each $\rho^{(t)}$ to the closest center $\rho^{*}_{\ell}$, and update the region of attraction $\mathcal{N}_{\ell}$
     if any region of attraction $\mathcal{N}_{\ell}$ is empty then
         Replace $\rho^{*}_{\ell}$ with a distant partition $\rho^{(t)}$ and update $\mathcal{N}_{\ell}=\{\rho^{(t)}\}$
     VI-search step:
     for $\ell$ in $1,\ldots,L$ do
         Update centers: $\rho^{*}_{\ell}\leftarrow\textsc{minVI}(\mathcal{N}_{\ell})$
         Update centers’ expected-VI: $l_{\ell}\leftarrow\textsc{EVI}(\rho^{*}_{\ell},\mathcal{N}_{\ell})$
         Update weights $w_{\ell}\leftarrow|\mathcal{N}_{\ell}|/T$
     Update loss $W\leftarrow\sum_{\ell=1}^{L}w_{\ell}\cdot l_{\ell}$
until change of $W$ is less than $\epsilon$.
Return $\{\rho^{*}_{1},\ldots,\rho^{*}_{L}\},\mathbf{w}=(w_{1}\ldots,w_{L})$
The search algorithm to find the $L$ centers that support the approximate WASABI posterior has a similar structure to the k-medoids algorithm. Essentially, given an initialization of the centers, the algorithm alternates between what we refer to as the N-update step, where the regions of attraction are updated, and the VI-search step, where the new center for each region of attraction corresponds to the minVI of each group.
More precisely, the N-update step assigns all points (the MCMC samples) to their closest center; this requires computing the VI distance between each point and each center. In the rare case when a partition is equally distant to two or more centers, it is randomly allocated among them. It is then followed by an optional outlier-search step, which is attempted when one region of attraction contains only one partition, meaning that the current center is not the closest to any other partition; this step tries to swap the current center with a sampled “outlier” partition, i.e. a partition that is far (in VI sense) from all centers, and checks if the objective function improves. Computing the VI distance between all points and centers is the most expensive part of this step, with complexity $O(nTL)$.
Next, the VI-search updates the center of each group given their new region of attraction, but unlike k-medoids, which restricts the search to the points within the group, we make use of more advanced algorithms to minimize the posterior expected VI, such as SALSO (Dahl et al., 2022). The complexity of this step depends on the algorithm selected, e.g. for SALSO (the default), the cost is $O(nTK_{\text{max}}^{2})$, where $K_{\text{max}}$ is the maximum number of clusters in the MCMC samples.
The algorithm iterates between these steps until a maximum number of iterations is reached, or the improvement in the Wasserstein distance is less than a given threshold $\epsilon$. Note that, since the VI distance is bounded above by $\log_{2}(n)$, we recommend choosing a threshold that depends on this quantity, such as $\epsilon=0.0001*\log_{2}(n)$. The basic algorithm is outlined in Algorithm 1, and a more detailed version can be found in Section B of the Appendix. As the algorithm only guarantees convergence to a local optimum, it is recommended to run it multiple times with different starting conditions. We consider several initialization strategies, including using the top $L$ partitions found with hierarchical clustering based on the PSM, or a strategy mimicking the k-means++ initialization (Vassilvitskii and Arthur, 2006), which randomly initializes the centers by selecting more spread out partitions among the samples. Further initialization details are provided in the Appendix (Section B). To reduce computational complexity, strategies employed in k-means can be adapted; for example, we allow a “mini-batch” option, where at each iteration, we consider a subsample of the MCMC partitions of size $B$. This reduces the cost to $O(nB(L+K_{\text{max}}^{2}))$ but is followed by a small number of iterations using the full set of MCMC partitions, similar to the two-stage k-means of Salman et al. (2011).
Choosing the number of particles $L$
To approximate the posterior distribution with WASABI, the practitioner must select the number of particles $L$. As $L\rightarrow T$, the Wasserstein distance reduces to zero, but it becomes more and more difficult to describe all centers. To achieve a balance between parsimony (small $L$) and a good enough approximation (small Wasserstein distance), we propose running the algorithm for a range of values of $L$ (usually from 1 to 10), and plotting the Wasserstein distance as a function of $L$. This produces an “elbow plot”, and the selected value of $L$ is the smallest that corresponds to a significant reduction of the Wasserstein distance, before the curve flattens (or decreases less substantially). See e.g. the left panel in Figure 2(c).
3.2 Describing and visualizing the particles
Clustering is often used for exploratory data analysis, and visualization of the partitions is a crucial part of the analysis to uncover and better understand patterns in the data. Similarly, visualization of the obtained particles is essential to better understand the structure of the posterior distribution.
For the simple slightly-bimodal data (Example 1), comparison of the particles is quite straightforward. We now consider a two-dimensional extension of this simple example, and in the following, outline several techniques to describe and visualize the particles that summarize the posterior, illustrated on this two-dimensional example.
Example 2.
Consider a two-dimensional Gaussian mixture with four components, each with mean (±$(\pm m,\pm m)$ located in one of the four quadrants and diagonal covariance matrix with unit variance. Figure 2(a) displays the data colored by the partition used in the data-generating process (for $m=1.3$). MCMC samples are obtained by fitting a diagonal location-scale Dirichlet process mixture (DPM) using the BNPmix R package (Corradin et al., 2021). The minVI partition (Figure 2(b)) merges three of the four components, which does not reflect the axis-aligned elliptical clusters of the DPM model. This highlights that a single-point estimate can be misleading when the posterior is multimodal. In fact, the WASABI elbow plot (Figure 2(c)) suggests that multiple estimators are useful for describing the posterior. We now provide several approaches for visualizing the WASABI particles.
(a) Scatterplot of the data
(b) minVI partition
(c) WASABI elbow plot
Figure 2: Two-dimensional extension of the bimodal example. Data is generated from a Gaussian mixture with four components (Panel 2(a)). The minVI merges three components (Panel 2(b)). To choose the number of particles $L$ in WASABI, we construct an elbow plot (Panel 2(c)) which suggests $L=3$ particles achieves a balance between parsimony and minimizing the objective.
Weight and number of clusters for each particle
Determining the number of clusters is an important problem, and with WASABI, we can compare the number of clusters identified by the different particles. In several experiments, we have found that different particles often tend to have different numbers of clusters. Moreover, the weights associated with the particles convey their importance in summarizing the posterior MCMC samples, as they correspond to the proportion of samples that are closer to each given particle. If a particle has a weight of 0.7, it means that 70% of the MCMC samples are closer to this particle, and can then be summarized by it. See the left panel of Figure 1(d) for a representation of these quantities for Example 1.
Visualizing particles, with the data colored by cluster membership
The number of clusters provides limited information on the partition itself. By visualizing the partition corresponding to each particle, side by side, and depicting the data points colored by their cluster allocation, we can identify how the clusters change. For data that is more than 2-dimensional, transformations can be applied, such as principal component analysis.
Example 2 (continued).
Figure 3(a) displays the partitions recovered by each of the particles. The second particle detects three clusters, merging together the two top ‘true’ components, while the third one recovers again three clusters, merging together the two left ‘true’ components. Notice how the minVI partition is not among the particles and is obtained as product of the multimodality and uncertainty in merging either the two top components (particle 2) or the two left components (particle 3). Additionally, unlike Example 1, the particles are not nested within each other and could not be recovered using a hierarchical clustering method.
Comparing the particles globally: visualizing the meet
For a global comparison of the particles, we can examine their meet. In the space of partitions, which forms a partially ordered set under the relation of set containment, the meet (denoted $\wedge$) corresponds to the greatest lower bound of a collection of partitions. Specifically, the meet of a set of partitions is the coarsest partition such that each of its clusters is fully contained within a cluster of every partition in the set. The meet’s clusters highlight which data points are always clustered together across all of the partitions from the set. Therefore, by visualizing the particles’ meet we can better identify parts of the data points that are consistently clustered together in WASABI.
Example 2 (continued).
Figure 3(c) (top) shows the cluster allocation of the meet partition of the particles. The four largest clusters recover the four quadrants that correspond to the four components used to generate the data, while the three remaining clusters corresponds to a small number of data points whose cluster assignment changes in particles 2 and 3. For example, the data points in meet cluster 5 were allocated in cluster 1 in particle 2 and in cluster 3 in particle 3.
(a) Particles’ cluster assignment.
(b) Posterior similarity matrix (PSM) for each particle’s region of attraction.
(c) Cluster allocation and EVIC for the particles’ meet.
(d) Collapsed WASABI similarity matrix.
(e) VIC between particle 2 and particle 3.
Figure 3: Several possible visualizations of the WASABI summary for Example 2: (a) particle’s scatterplot with cluster assignment, (b) posterior similarity matrix for the samples within each region of attraction, (c) the particle’s meet, (d) posterior similarity matrix using the WASABI approximation collapsed to meet’s clusters, (e) comparison of two particles using VI contribution.
Comparing pairs of particles: visualizing the VI-contribution
Sometimes, the visualization of particles side by side does not allow for a full understanding of the differences in the cluster assignment, especially when the number of data points is large. Instead, we can highlight the differences between two particles by considering each point’s contribution to the VI distance between those particles.
Definition 3 (VI contribution).
The VI distance between two partitions $\rho_{1}$ and $\rho_{2}$ can be written as the sum over all data points of a non-negative contribution, $\text{VI}(\rho_{1},\rho_{2})=\sum_{i=1}^{n}\text{VIC}_{i}(\rho_{1},\rho_{2})$, where each point’s contribution to the VI is defined as
$\text{VIC}_{i}(\rho_{1},\rho_{2})$ $=\frac{1}{n}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c_{1,i}=c_{1,i^{\prime}})}{n}\right)+\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{2,i}=c_{2,i^{\prime}})}{n}\right)\right.$
$\quad\left.-2\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{1,i}=c_{1,i^{\prime}},c_{2,i}=c_{2,i^{\prime}})}{n}\right)\right].$
Note first that if two points are clustered together under the meet $\rho_{1}\wedge\rho_{2}$, then their VI contribution will be the same. Moreover, the VI contribution will be zero for a point only if it is clustered with the same set of points in both partitions. More precisely, if a cluster $C_{h}\in\rho_{1}$ and $C_{h}\in\rho_{2}$ (and consequently $C_{h}\in\rho_{1}\wedge\rho_{2}$), then $\text{VIC}_{i}(\rho_{1},\rho_{2})=0$ for any $i\in C_{h}$. At the other extreme, if a data point is clustered with a completely different set of data points, then it’s VI contribution will be large, especially if the respective clusters are large, i.e. if $i\in C_{h_{1}}$ in $\rho_{1}$ and $i\in C_{h_{2}}$ in $\rho_{2}$, and $|C_{h_{1}}\cap C_{h_{2}}|=1$, then $\text{VIC}_{i}(\rho_{1},\rho_{2})=(\log_{2}(n_{h_{1}})+\log_{2}(n_{h_{2}}))/n$. A full derivation of the VI contribution is included in Appendix C.1.
Example 2 (continued).
Coloring the data points based on their VI-contribution can help pinpoint areas within the dataset where the two particles differ, enabling more detailed investigation. Figure 3(e) compares particles 2 and 3, with the lowest VI contribution for the bottom right quadrant, highlighting that these points are mostly clustered together, and the highest for points on the border, suggesting more drastic changes in who they are clustered with. See also the Appendix, for an alternative visualization.
Since the VI contribution is constant for all points that belong to the same meet cluster, we can define the “VI contribution by group” (VICG), i.e. the aggregate VIC for all points belonging to the same meet cluster.
Definition 4 (VI contribution by group).
The VI distance between the two partitions can be written as the sum over all clusters in the meet of a non-negative “group contribution”, $\text{VI}(\rho_{1},\rho_{2})=\sum_{k_{1}=1}^{K_{1}}\sum_{k_{2}=1}^{K_{2}}\text{VICG}_{k_{1},k_{2}}(\rho_{1},\rho_{2})$, where
$\text{VICG}_{k}(\rho_{1},\rho_{2})$ $=\frac{n_{k_{1},k_{2}}}{n}\left[\log_{2}\left(\frac{n_{k_{1},+}}{n}\right)+\log_{2}\left(\frac{n_{+,k_{2}}}{n}\right)-2\log_{2}\left(\frac{n_{k_{1},k_{2}}}{n}\right)\right].$
The “groups” of the VI contribution by group correspond to the meet’s clusters, which are also found as the pairwise intersection of each of the particles’ clusters. While inspecting the VIC for each data point highlights the subtle differences between two partitions, by plotting the VICG we can easily highlight high-level changes in the two partitions. In fact, major changes in the cluster assignment, such as a large cluster being split into two equally-sized subclusters, would have a significant contribution to the VI, but if the number of points in those clusters is large, the individual contribution may be negligible, since each point is still clustered together with many of the same points. However, by aggregating the contribution across all points in the meet’s cluster in the VICG, these large-scale differences become more visible. Figure 10 in the Appendix displays the VICG for the data in Example 2.
Investigating the co-clustering properties of the meet’s clusters
When the number of clusters in the particles’ meet is large, we can benefit from a better understanding of how often these clusters are combined together across the particles. The proposed visualization achieves this, while also comparing the meet with each of the particles. Specifically, for any pair of meet clusters $C_{k},C_{h}\in\bigwedge_{\ell=1}^{L}\rho_{\ell}^{*}$, we consider their WASABI posterior similarity, defined as the posterior probability that any two points in these meet clusters, $i\in C_{k}$ and $i^{\prime}\in C_{h}$, are clustered together as approximated by WASABI, $\text{Pr}(c_{i}=c_{i^{\prime}}|\bm{y})\approx\sum_{\ell=1}^{L}w_{\ell}\mathbbm{1}(\{C_{k}\cup C_{h}\}\in\rho^{*}_{\ell})$. In other words, we measure how often the meet’s clusters are clustered together in the particles using their weights. This is equivalent to using WASABI to approximate the posterior similarity matrix, grouped by the meet’s clusters (instead of individual data points, since it will be equal among points with the same allocation in the meet). We can visualize this PSM by grouping together the meet’s clusters that form each of the clusters in a particle, and thus providing further investigation in the particle’s structure. See for example Figure 3(d) for a comparison of particles 2 and 3 with the meet, for Example 2. The meet’s clusters are aggregated within the particle’s clusters in the rows, with labels shown on the left margin.
Determining the uncertainty in each region of attraction: inspecting each particle’s posterior similarity matrices
The WASABI particles summarizes the posterior distribution with several “point estimates” that can be visualized for easy understanding of the posterior structure. Uncertainty remains around each of the particles, as they are the representative partition for a group of MCMC samples (their “region of attraction”). To understand the uncertainty in each region of attraction, we can visualize the PSM of each region, i.e. calculating it only using the MCMC samples that are closest to the corresponding particle. For example, Figure 3(b) displays the PSM for each of the particles’ regions of attraction for Example 2, using heatmaps by grouping (and labeling) each point into their cluster assignment. Additionally, for each of the particles we can compute the (normalized) Expected VI distance, where the expectation is taken only with respect to the MCMC samples in their region of attraction. This quantifies the spread of the samples around the particle; for example, there is more variability in the first region of attraction (characterized by single, large cluster) compared to the second (see plot titles in Figure 3(b)). An alternative visualization that groups data points by the meet’s cluster allocations is provided in Figure 8 of the Appendix.
Determining the uncertainty in each point’s cluster allocation: the contribution to the expected-VI
While the VIC enables the comparison of two specific partitions, the same decomposition of the VI can also be applied in the context of the expected VI to obtain a measure of uncertainty in each point’s cluster allocation. Essentially, the expected VI quantifies the variability of the posterior around a particular partition. By analyzing how each point contributes to the EVI, we can derive a measure of uncertainty for the cluster assignment of each point.
Let $\rho^{*}$ be an estimator for a partition (such as the minVI, or one of the WASABI particles). We define the contribution of the $i$th data point to the expected VI of $\rho^{*}$ in Definition 5.
Definition 5 (EVI contribution).
The contribution of the $i$th data point to the EVI of $\rho^{*}$ is
$\text{EVIC}_{i}(\rho^{*})$ $=\frac{1}{n}\left\{\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c^{*}_{i}=c^{*}_{i^{\prime}})}{n}\right)+\mathbb{E}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{i}=c_{i^{\prime}})}{n}\right)\mid\bm{y}\right]\right.$
$\quad\left.-2\mathbb{E}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{i}=c^{*}_{i^{\prime}},c_{i}=c_{i^{\prime}})}{n}\right)\mid\bm{y}\right]\right\},$
where the expectation is taken with respect to $\pi(\rho\mid\bm{y})$, the posterior distribution over the space of clusterings. The EVI of $\rho^{*}$ is:
$\text{EVI}(\rho^{*})=\sum_{i=1}^{n}\text{EVIC}_{i}(\rho^{*}).$
Note that computing the contribution of the $i$th data point the EVI of $\rho^{*}$ requires computing the expectation with respect to the posterior over the space of partitions. This can be approximated either based on the MCMC draws:
$\widehat{\text{EVIC}}_{i}(\rho^{*})$ $=\frac{1}{n}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c^{*}_{i}=c^{*}_{i^{\prime}})}{n}\right)+\frac{1}{T}\sum_{t=1}^{T}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{(t)}_{i}=c^{(t)}_{i^{\prime}})}{n}\right)\right.$
$\quad\left.-2\frac{1}{T}\sum_{t=1}^{T}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{i}=c^{*}_{i^{\prime}},c^{(t)}_{i}=c^{(t)}_{i^{\prime}})}{n}\right)\right],$
or based on the WASABI posterior:
$\widehat{\widehat{\text{EVIC}}}_{i}(\rho^{*})$ $=\frac{1}{n}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c^{*}_{i}=c^{*}_{i^{\prime}})}{n}\right)+\sum_{l=1}^{L}w_{l}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{l,i}=c^{*}_{l,i^{\prime}})}{n}\right)\right.$
$\quad\left.-2\sum_{l=1}^{L}w_{l}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{i}=c^{*}_{i^{\prime}},c^{*}_{l,i}=c^{*}_{l,i^{\prime}})}{n}\right)\right].$
Approximating the contribution of the $i$th data point to the EVI based on the WASABI posterior has a clear computational advantage ($O(nL)$ as opposed to $O(nT)$ with $L\ll T$); in addition, note that when $\rho^{*}$ is one of the particles, every data point in the same cluster of the meet will have the same contribution to the VI, further reducing computations and enhancing interpretation. For example, clusters in the meet with a WASABI approximated EVI of zero, are always clustered together and never with any other data points, across all particles. Lastly, we note that the EVI of the $l$th particle, $\rho^{*}_{l}$ can also be taken with respect to the posterior restricted to its region of attraction to shed light on which points are more uncertain (or stable) in their cluster allocation within the region of attraction.
Example 2 (continued).
The contribution to the expected VI for the meet of the particles is displayed in Figure 3(c) (bottom). By analyzing the patterns in this figure, we notice that the points with the largest EVIC are the ones at the border between the four large clusters (depicted in Figure 3(c) (top) and represented here with different shapes). Moreover, the points in the top-left component (labeled with 1) have a slightly larger EVIC than the points in the bottom-right component (labeled with 6), reflecting that their cluster allocation changes more across the MCMC samples, while the cluster assignment of the points in cluster 6 change less.
4 Experiments
We now consider several synthetically generated datasets to illustrate scenarios where WASABI can be beneficial in explaining the uncertainty in the posterior distribution. We consider two main scenarios where the posterior distribution displays high uncertainty: (1) when the model is correctly specified, but the mixture components are not well-separated, and (2) when the model is misspecified.
First, we consider scenario (1), by examining different levels of cluster separation in Example 2. Specifically, several values of $m$, defining the means (±$(\pm m,\pm m)$ of the four mixture components, are considered (see Figure 11 in the Appendix). For each value of $m$, we generate 10 datasets. For small values of $m$, the four mixture components overlap, making the distribution concentrated around the partition with one cluster. In contrast, when $m$ is large, the four components are well separated and easy to distinguish.
To quantify the uncertainty in the posterior, we can measure the posterior distribution’s spread around its center by the EVI. This is displayed in the left panel of Figure 4, showing how the posterior uncertainty starts low and increases when there is a moderate level of cluster separation, and then decreases when the cluster separation is high. The right panel of Figure 4 instead shows the improvement in the Wasserstein distance achieved by WASABI, quantified by $\textrm{EVI}(\rho^{*})-W_{\text{VI}}(\hat{\pi},q^{*})$, when summarizing with $L=10$ particles instead of a single estimator. These figures show that WASABI can explain the variation in the posterior the most when the posterior uncertainty is highest (i.e. for values of $m\approx 1.3$).
Figure 4: Experiment with different levels of cluster separation, determined by $m$. Left panel: posterior spread around the minVI estimator, quantified by its expected VI. Right panel: improvement in Wasserstein distance achieved by WASABI for different values of $m$; the points are colored by the number of clusters in the corresponding minVI estimate.
In scenario (2), we explore several examples of misspecified models with data generated from different distributions: (a) a bivariate Gaussian distribution truncated to the unit square, and (b) a bivariate skewed-t distribution (assuming independence). To investigate the effect of misspecification, in (a), the standard deviation is varied, so that small values are well-specified and larger values resemble a uniform distribution on the unit square; and in (b), the degrees of freedom and skewness parameter are varied. A detailed description of these synthetic experiments can be found in Section D of the Appendix.
Figure 5: Synthetic experiment with misspecified models. The top two rows show results for truncated Gaussian and skewed-t ($df=5$) distributions. Left: improvement in Wasserstein distance by WASABI versus posterior spread (EVI); central/right: particle scatterplots with cluster assignments and a summary plot, for one synthetic dataset. The bottom row shows Wasserstein improvement for skewed-t with varying degrees of freedom.
Figure 5 displays the improvement in Wasserstein distance achieved by WASABI as a function of the posterior spread (measured by EVI) for (a) the truncated Gaussian distribution (top row, left panel), and (b) for the skewed-t distribution with $df=5$ (second row, left panel). As expected, overall WASABI tends to explain higher variations in the posterior when the EVI is larger. This also correlates with increasing misspecification; for the truncated Gaussian distribution (a), the EVI increases with the standard deviation, and for the skewed-t distribution with $df=5$ (b), larger values of the skewness parameter result in higher EVI and larger improvement by WASABI. The same improvement is visible in the bottom row (all panels) for the skewed-t distribution with varying degrees of freedom, where we can also observe that most of the variation is driven by the skewness parameter, but for smaller degrees of freedom, higher posterior uncertainty (and improvement achieved by WASABI) and observed. Finally, the central and right panels in the first two rows provide a summary of the data and particles for one representative example, displaying median values of EVI and improvement in Wasserstein distance; for the truncated Gaussian, this corresponds to $\sigma=0.4$, while for the skewed-t distribution it corresponds to $df=5$ and skewness parameter equal 2.
5 Real data analysis
The human papillomavirus (HPV) vaccine, a key tool in reducing cervical and other cancers, is routinely offered to 14-year-old girls in Scotland. However, decline in uptake is a global concern, underscoring the need to examine factors that influence this pattern. Here, we focus on studying how neighborhood deprivation may affect HPV vaccine uptake in Scotland. The data [^2] Data available through Public Health Scotland: https://scotland.shinyapps.io/ScotPHO_profiles_tool/ consists of the percentage of HPV uptake, denoted by $y_{i}$, and the deprivation level, denoted by $x_{i}$ and measured as the proportion of young people living in the most income deprived quintile, across the $n=1279$ Scottish Intermediate Zones. For a visualization of HPV uptake across these zones, see the left panel of Figure 6.
Simple linear regression models only allow us to understand how deprivation affects mean uptake. Importantly, we are also interested in the left tail of relatively low uptake and how it is influenced by deprivation. This motivates the use of models for density regression, which allow the entire density of HPV uptake to change as a function of deprivation. We focus on Bayesian dependent mixtures (Wade and Inácio, 2025), which allow flexible density regression and are validated by well-established theory (Ghosal and Van der Vaart, 2017). In particular, we employ a Bayesian mixture of linear regressions (Rodríguez-Álvarez et al., 2025), assuming $Y_{i}|x_{i}\sim\sum_{k=1}^{\infty}w_{k}N(\alpha_{k}+\beta_{k}x_{i},\sigma_{k}^{2})$, where the weights have a stick-breaking construction $w_{k}=v_{k}\prod_{j<k}(1-v_{j})$ and $v_{k}\sim\text{Beta}(1,\alpha)$. MCMC is used to approximate the posterior distribution and estimate the conditional densities of vaccine uptake given different levels of neighborhood deprivation (center panel of Figure 6). There is a tail of regions with less uptake, which is more present as deprivation increases. To better understand this behavior and the underlying regions, WASABI can be used to summarize the MCMC draws of clusterings associated to the Bayesian mixture model.
Figure 6: HPV vaccine uptake in Scotland. Left: Visualization of HPV uptake rates across neighborhoods. Center: Estimated conditional density (with pointwise credible intervals) of HPV vaccine uptake by neighborhood deprivation, showing lower uptake and a bimodal pattern for higher deprivation neighborhoods. Right: Map of the second WASABI particle’s cluster assignments.
The posterior on the clustering structure reflects groups of neighborhoods with similar relation between uptake and deprivation and overall level of uptake. To summarize and understand this posterior, WASABI finds two particles with weights $w_{1}=0.87$ and $w_{2}=0.13$; the first particle corresponds to the partition with one cluster, while the second one finds two clusters. This reflects the possibility of two groups of regions with different patterns, although the separation is not clear. The two clusters of the second particle (depicted in the right panel of Figure 6) are characterized by the two linear regression equations, respectively $(\widehat{\alpha}_{1},\widehat{\beta}_{1})=(90.05,-0.09)$ for the higher uptake cluster and $(\widehat{\alpha}_{2},\widehat{\beta}_{2})=(76.18,-0.12)$ for the lower uptake cluster, whose lower uptake becomes even more pronounced as deprivation decreases. Interestingly, from the map of the second particle (right panel of Figure 6), we can see a spatial correlation in the cluster assignments, and that surprisingly the two main cities in Scotland, Edinburgh and Glasgow, have very different patterns. Edinburgh seems to be comprised mostly of lower-uptake regions, while Glasgow of higher-uptake regions. Further investigations are required to understand this difference, possibly caused by factors, such as high school catchments, the effect of word-of-mouth, and local advertisement and health facilities.
6 Discussion
Understanding the uncertainty characterized by the posterior distribution is challenging in complex, high-dimensional Bayesian models. Additionally, as the posterior is intractable in such settings, algorithms for approximate inference, such as MCMC, are employed, which can return tens to hundreds of thousands of approximate posterior draws. In this work, we focus on clustering, where the posterior is defined on the high-dimensional, partially-ordered space of partitions. In this non-standard setting, we propose a WASserstein Approximation for Bayesian cluserIng (WASABI) to summarize the posterior draws with a small number of weighted partition estimates. Through extensive simulations and analyses, we showcase its utility in practical settings when some clusters are not well separated or the model may be misspecified. Indeed, while many applications focus on a single clustering estimate, we demonstrate how it can be misleading in such settings and how WASABI captures the different modes of clustering.
The Wasserstein approximation requires selecting a metric on the partition space. We focus on the variation of information (Meilă, 2007), also exploiting its properties to propose novel tools to assess uncertainty in the cluster assignment of individual points. However, this can be easily extended to other choices, such as Binder’s loss, ARI, or generalized versions (Dahl et al., 2022), or to more general settings, e.g. Buch et al. (2024) develop a Bayesian approach to clustering in the presence of noise.
An efficient algorithm, along with various tools to visualize and quantify uncertainty, is provided in the R package at https://github.com/cecilia-balocchi/WASABI. In practice, only a small number of iterations are typically required, with each one scaling linearly with the number of data points and MCMC samples. To reduce the cost of the latter, WASABI allows subsampling of the MCMC draws. For the former, future work will explore combining WASABI with distributed methods for Bayesian clustering (Ni et al., 2020).
WASABI is useful for improved understanding of the posterior, but could also be relevant for posterior compression in order to reduce both storage (saving only the WASABI particles and weights) and the computational cost of predictive inference. For example, density estimation can be expensive, depending on the size of the grid or number of new locations. For Example 1, density estimates are provided in Section D.1 of the Appendix, where the WASABI-based estimates significantly reduce the cost by a factor of $1,000$ compared to MCMC, at only a small increase in error.
Beyond clustering, the Wasserstein approximation could be beneficial for other Bayesian models. Future work, will investigate settings, such as variable selection, trees, networks, and more.
Appendix
Appendix A Proof to Proposition 1
Proof.
Let $\bm{w}^{*}=(w^{*}_{1},\ldots,w^{*}_{L})$ and $\bm{\rho}^{*}=\{\rho^{*}_{1},\ldots,\rho^{*}_{L}\}$ respectively denote the vector of weights and the set of atoms of $q^{*}(\cdot)=\sum_{\ell=1}^{L}w^{*}_{\ell}\delta_{\rho^{*}_{\ell}}(\cdot)$. To solve (2) we need to find the optimal $\bm{w}^{*}$ and $\bm{\rho}^{*}$, but also for a given $q$ we need to find the optimal coupling $J^{*}(q)$ that minimizes $\sum_{\rho,\rho^{\prime}}d_{VI}(\rho,\rho^{\prime})J(\rho,\rho^{\prime})$ and that satisfies the marginal constraints.
Note that we can equivalently frame the problem as finding the optimal set of atoms $\bm{\rho}^{*}$ and the optimal $J^{*}$ such that $\sum_{\rho}\sum_{\ell=1}^{L}d_{VI}(\rho,\rho^{*}_{\ell})J^{*}(\rho,\rho^{*}_{\ell})$ is minimized and $\sum_{\ell=1}^{L}J^{*}(\rho,\rho^{*}_{\ell})=\pi(\rho|\bm{y})$. From $J^{*}$ we can recover the set of weights $\bm{w}^{*}$ as $w^{*}_{\ell}=\sum_{\rho}J^{*}(\rho,\rho^{*}_{\ell})$ for $\ell=1,\ldots,L$.
This follows from the fact that, for any $J$ and for each $\rho$, $\rho^{*}_{\ell}$, we have that $J(\rho,\rho^{*}_{\ell})\leq\pi(\rho|\bm{y})$; this means that $w^{*}_{\ell}=\sum_{\rho}J^{*}(\rho,\rho^{*}_{\ell})\leq\sum_{\rho}\pi(\rho|\bm{y})\leq 1$ and that $\sum_{\ell=1}^{L}w^{*}_{\ell}=\sum_{\ell=1}^{L}\sum_{\rho}J^{*}(\rho,\rho^{*}_{\ell})=\sum_{\rho}\sum_{\ell=1}^{L}J^{*}(\rho,\rho^{*}_{\ell})=\sum_{\rho}\pi(\rho|\bm{y})=1$.
To find the optimal $J^{*}$ given $\bm{\rho}^{*}$, note that we can find the values $J^{*}(\rho,\rho^{*}_{\ell})$ independently for each $\rho$. Thus for each $\rho$ we can recover these values by minimizing $\sum_{\ell=1}^{L}d_{VI}(\rho,\rho_{\ell})J^{*}(\rho,\rho_{\ell})$ over $(J^{*}(\rho,\rho^{*}_{\ell}))_{\ell=1}^{L}$ with the constraint that $\sum_{\ell=1}^{L}J^{*}(\rho,\rho^{*}_{\ell})=\pi(\rho|\bm{y})$. This can be recognized as a linear program, which is solved by $J^{*}(\rho,\rho^{*}_{\ell})=\pi(\rho|\bm{y})$ for $\ell=\ell^{*}(\rho)$, where $\ell^{*}(\rho)=\operatornamewithlimits{argmin}_{\ell}d_{VI}(\rho,\rho_{\ell})$, and $J^{*}(\rho,\rho^{*}_{\ell})=0$ for $\ell\neq\ell^{*}(\rho)$. Note that in the case where there are multiple $\rho^{*}_{\ell}$ minimizing the VI distance with $\rho$, then a continuum of solutions exists; we focus on the extreme cases of allocating $\rho$ to all equally distant particles and consider all possible solutions. This can be done in practice by randomly allocating $\rho$ to one of the closest particles $\rho^{*}_{\ell}$, and setting $J^{*}(\rho,\rho^{*}_{\ell})=\pi(\rho|\bm{y})$. ∎
Appendix B Detailed algorithm and initialization strategies
The algorithm to find the approximate WASABI posterior (see Section 3.1) is akin to the k-medoids approach and only guarantees convergence to a local optimum. Therefore, it is standard practice to run the algorithm multiple times with different initializations, to increase the chance of finding a better (lower-loss) solution. Below, we describe in detail the initialization strategies implemented:
  • Average Linkage Initialization: A hierarchical clustering is performed on the posterior similarity matrix using average linkage, and the $L$ resulting partitions with number of clusters ranging from $1$ to $L$ (from appropriate tree cuts) are selected as initial centers. This approach specifies the initialization with diverse partitions that represent stable co-clustering patterns in the MCMC sample.
  • Complete Linkage Initialization: Similarly, hierarchical clustering using complete linkage is applied to the posterior similarity matrix. The dendrogram is cut at levels corresponding to $1$ up to $L$ clusters, and the resulting $L$ partitions are used as the initial medoids. Complete linkage generally favors compact, small-diameter clusters.
  • Fixed Initialization: This strategy allows the user to directly specify a set of $L$ fixed partitions to be used as initial centers. This is useful, for example, if the user wishes to include certain known clusterings (such as the minVI estimator, MAP clustering, or expert-proposed solutions) as initial candidates. Alternatively, the EVI can be computed for all MCMC samples, and the $L$ partitions with the smallest values are passed as starting points.
  • k-means++ style Initialization (+++++ +): Analogous to the k-means++ technique (Vassilvitskii and Arthur, 2006), this strategy promotes diversity among initial centers by iteratively choosing the next center with probability proportional to its VI distance from the closest already chosen center. This increases the likelihood that the $L$ initial medoids will be well separated in partition space, which often leads to better local optima.
  • Random Partition Initialization: In this approach, each data point is randomly assigned to one of $L$ groups, and the center for each group is chosen based on these assignments. This tends to produce initial medoids that are close together in partition space, often near the “center” of the sample distribution. Although not suited for finding very distinct particles, it may be useful for robustness checks.
  • Enhanced Initialization (+++++++ + +): This extends the +++++ + initialization by running the +++++ + initialization search on the set of the MCMC samples combined with the partitions obtained from the average and complete linkage strategies. The intention is to expand the diversity of the initial set and include particularly informative candidates, possibly leading to a broader exploration of the solution space.
  • Hierarchical Super-Initialization (topvi): In this method, the expected VI is computed for all the partitions generated by the average linkage, complete linkage, and (if supplied) fixed initialization methods. The $L$ best partitions among these (with lower EVI) are then selected as initial centers. This approach mimics the idea of selecting the partitions in the MCMC sample with lowest EVI, while avoiding the computational complexity of computing the EVI for all MCMC samples.
Note: When practical, it is helpful to combine several strategies and perform multiple runs with each, then select the solution yielding the lowest objective (Wasserstein) distance. In particular, using k-means++ or hierarchical strategies offers a good trade-off between computational cost and solution diversity, while random initialization is useful for unbiased robustness assessment.
Algorithm 2 WASABI: Finding the approximate WASABI posterior (detailed steps)
Input: MCMC samples $\{\rho^{(t)}\}_{t=1}^{T}$, number of particles $L$, initialization method init, tolerance $\epsilon$, maximum number of iterations max_iter
Initialize $\rho^{*}_{1:L}$ using method init
$i\leftarrow 0$
repeat
     $i\leftarrow i+1$
     N-update step:
     Compute $\text{VI}(\rho^{(t)},\rho^{*}_{\ell})$ for all $t$ and $\ell$
     Assign each $\rho^{(t)}$ to the closest center $\rho^{*}_{\ell}$, and update the region of attraction $\mathcal{N}_{\ell}$
     if any region of attraction $|\mathcal{N}_{\ell}|=0$ then $\triangleright$ Empty region of attraction
         For such $\ell$, compute $p_{\ell,t}\propto\text{VI}(\rho^{(t)},\rho^{*}_{\ell})$
         Sample $t^{*}$ from $1,\ldots,T$ according to $p_{\ell,t}$
         Set $\rho^{*}_{\ell}\leftarrow\rho^{(t^{*})}$ and $\mathcal{N}_{\ell}\leftarrow\{t^{*}\}$
     Outlier-check step:
     if any region of attraction $|\mathcal{N}_{\ell}|=1$ then
         For such $\ell$, compute $p_{\ell,t}\propto\min_{\ell^{\prime}}\text{VI}(\rho^{(t)},\rho^{*}_{\ell^{\prime}})$
         Sample $t^{*}$ from $1,\ldots,T$ according to $p_{\ell,t}$
         if objective decreases upon replacement then
              Set $\rho^{*}_{\ell}\leftarrow\rho^{(t^{*})}$ and $\mathcal{N}_{\ell}\leftarrow\{t^{*}\}$
     VI-search step:
     for $\ell$ in $1,\ldots,L$ do
         Update centers: $\rho^{*}_{\ell}\leftarrow\textsc{minVI}(\mathcal{N}_{\ell})$triangleright Different search-algorithms can be used
         Update centers’ expected-VI: $l_{\ell}\leftarrow\textsc{EVI}(\rho^{*}_{\ell},\mathcal{N}_{\ell})$
         Update weights $w_{\ell}\leftarrow|\mathcal{N}_{\ell}|/T$
     Update loss $W\leftarrow\sum_{\ell=1}^{L}w_{\ell}\cdot l_{\ell}$
[…]
Return $\{\rho^{*}_{1},\ldots,\rho^{*}_{L}\},\mathbf{w}=(w_{1}\ldots,w_{L})$
Appendix C Additional material on describing and visualizing the particles
In Section 3.2, we introduce Example 2, where the data is generated from four Gaussian mixtures. We fit a diagonal location-scale Dirichlet Process Mixture model, using the R package BNPmix (Corradin et al., 2021), run for 1,000 iterations after a burn-in of 10,000 iterations. In the posterior similarity matrix shown in Figure 7, uncertainty about the number of clusters is apparent, but a subtle pattern of four equal-sized blocks can be observed, even though some of these seem to be often merged together.
Figure 7: Posterior similarity matrix for the two-dimensional Gaussian mixture with four components (Example 2).
Visualization of the particles’ posterior similarity matrices
Figure 8 displays the posterior similarity matrices (PSMs) for the regions of attraction associated with each WASABI particle in Example 2. While Figure 7 (in Section 3.2) presents PSMs labeled by each particle’s own clusters (providing insight into the uncertainty structure around each specific clustering) here, the rows and columns are labeled according to the clusters defined by the meet of all particles. This common labeling facilitates direct comparison of cluster uncertainty across different regions of attraction, helping to identify points or groups whose allocation is most variable or most consistent between the particles represented by the WASABI summary.
Figure 8: Posterior similarity matrices for each region of attraction corresponding to the WASABI particles in Example 2. In each heatmap, rows and columns are labeled according to the meet’s clusters, enabling direct comparison of the PSM structure across different regions of attraction.
C.1 Contribution to the VI
To gain an understanding of the uncertainty in the allocation of each data point in a particular clustering, we compute each data point’s contribution to the expected VI. This can be computed by first recalling that the VI Meilă (2007) between two clusterings $\rho_{1}$ and $\rho_{2}$ is defined as:
$\text{VI}(\rho_{1},\rho_{2})$ $=\text{H}(\rho_{1})+\text{H}(\rho_{2})-2\text{MI}(\rho_{1},\rho_{2})$
$=-\sum_{j_{1}=1}^{K_{1}}\frac{n_{j_{1},+}}{n}\log_{2}\left(\frac{n_{j_{1},+}}{n}\right)-\sum_{j_{2}=1}^{K_{2}}\frac{n_{+,j_{2}}}{n}\log_{2}\left(\frac{n_{+,j_{2}}}{n}\right)-2\sum_{j_{1}=1}^{K_{1}}\sum_{j_{2}=1}^{K_{2}}\frac{n_{j_{1},j_{2}}}{n}\log_{2}\left(\frac{n_{j_{1},j_{2}}n}{n_{j_{1},+}n_{+,j_{2}}}\right),$
where the first two terms represent the entropy of the two clusterings and the last term is the mutual information between the clusterings. The counts $n_{j_{1},j_{2}}$ represent the cross-tabulation between the clusterings, i.e. $n_{j_{1},j_{2}}$ is the number of data points in cluster $j_{1}$ in $\rho_{1}$ and cluster $j_{2}$ in $\rho_{2}$, and $K_{1}$ and $K_{2}$ are the number of clusters in $\rho_{1}$ and $\rho_{2}$, respectively. We can equivalently express the VI as:
$\text{VI}(\rho_{1},\rho_{2})$ $=\sum_{i=1}^{n}\frac{1}{n}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c_{1,i}=c_{1,i^{\prime}})}{n}\right)+\sum_{i=1}^{n}\frac{1}{n}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{2,i}=c_{2,i^{\prime}})}{n}\right)$
$\quad-2\sum_{i=1}^{n}\frac{1}{n}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{1,i}=c_{1,i^{\prime}},c_{2,i}=c_{2,i^{\prime}})}{n}\right),$
where $c_{1,i}$ and $c_{2,i}$ represent the allocation of the $i$th data point, for $i=1,\ldots,n$, in the clusterings $\rho_{1}$ and $\rho_{2}$ respectively. Therefore, we define the contribution of the $i$th data point to the VI between $\rho_{1}$ and $\rho_{2}$ in Definition 6.
Definition 6 (Contribution to the VI).
The contribution of the $i$th data point to the VI between the clusterings $\rho_{1}$ and $\rho_{2}$ is
$\text{VIC}_{i}(\rho_{1},\rho_{2})$ $=\frac{1}{n}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c_{1,i}=c_{1,i^{\prime}})}{n}\right)+\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{2,i}=c_{2,i^{\prime}})}{n}\right)\right.$
$\quad\left.-2\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{1,i}=c_{1,i^{\prime}},c_{2,i}=c_{2,i^{\prime}})}{n}\right)\right],$
and the VI distance is:
$\text{VI}(\rho_{1},\rho_{2})=\sum_{i=1}^{n}\text{VIC}_{i}(\rho_{1},\rho_{2}).$
When comparing two clusterings (e.g. two particles), this can be useful to understand which data points contribute most to the distance, as well as those with zero (or low) contribution, whose allocation is the same between the two clusterings. Note that a desirable property of the VI is that it is aligned with the lattice of partitions. Specifically, it is both vertically aligned in the Hasse diagram: if $\rho_{1}\geq\rho_{2}\geq\rho_{3}$, then the distance between $\rho_{1}$ and $\rho_{3}$ is the vertical sum of the distances between $\rho_{1}$ and $\rho_{2}$ and between $\rho_{2}$ and $\rho_{3}$, as well as horizontally aligned: the distance between any two partitions $\rho_{1}$ and $\rho_{2}$ is the horizontal sum of the distances between each partition and the meet of the two partitions. Thus, we have that the VI contribution will be the same for all data points belonging to same cluster in the meet $\rho_{1}\wedge\rho_{2}$; allowing for more efficient computation.
Further, we define the contribution of the $i$th data point to the expected VI of a clustering $\rho^{*}$ in Definition 7.
Definition 7 (Contribution to the EVI).
The contribution of the $i$th data point to the EVI of the clustering $\rho^{*}$ is
$\text{EVIC}_{i}(\rho^{*})$ $=\frac{1}{n}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c^{*}_{i}=c^{*}_{i^{\prime}})}{n}\right)+\mathbb{E}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c_{i}=c_{i^{\prime}})}{n}\right)\mid\bm{y}\right]\right.$
$\quad\left.-2\mathbb{E}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{i}=c^{*}_{i^{\prime}},c_{i}=c_{i^{\prime}})}{n}\right)\mid\bm{y}\right]\right],$
where the expectation is taken with respect to $\pi(\rho\mid\bm{y})$, the posterior distribution over the space of clusterings. The EVI of $\rho^{*}$ is:
$\text{EVI}(\rho^{*})=\sum_{i=1}^{n}\text{EVIC}_{i}(\rho^{*}).$
Remark (Approximating the contribution to the EVI).
Computing the contribution of the $i$th data point the EVI of $\rho^{*}$ requires computing the expectation with respect to the posterior over the space of partitions. This can be approximated either based on the MCMC draws:
$\widehat{\text{EVIC}}_{i}(\rho^{*})$ $=\frac{1}{n}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c^{*}_{i}=c^{*}_{i^{\prime}})}{n}\right)+\frac{1}{T}\sum_{t=1}^{T}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{(t)}_{i}=c^{(t)}_{i^{\prime}})}{n}\right)\right.$
$\quad\left.-2\frac{1}{T}\sum_{t=1}^{T}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{i}=c^{*}_{i^{\prime}},c^{(t)}_{i}=c^{(t)}_{i^{\prime}})}{n}\right)\right],$
or based on the WASABI posterior:
$\widehat{\widehat{\text{EVIC}}}_{i}(\rho^{*})$ $=\frac{1}{n}\left[\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{\mathbf{1}(c^{*}_{i}=c^{*}_{i^{\prime}})}{n}\right)+\sum_{l=1}^{L}w_{l}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{l,i}=c^{*}_{l,i^{\prime}})}{n}\right)\right.$
$\quad\left.-2\sum_{l=1}^{L}w_{l}\log_{2}\left(\sum_{i^{\prime}=1}^{n}\frac{1(c^{*}_{i}=c^{*}_{i^{\prime}},c^{*}_{l,i}=c^{*}_{l,i^{\prime}})}{n}\right)\right].$
Approximating the contribution of the $i$th data point to the EVI based on the WASABI posterior has a clear computational advantage ($O(nL)$ as opposed to $O(nT)$ with $L\ll T$); in addition, note that when $\rho^{*}$ is one of the particles, every data point in the same cluster of the meet will have the same contribution to the VI, further reducing computations and enhancing interpretation. For example, clusters in the meet with a Wasserstein-VI approximated EVI of zero, are always clustered together and never with any other data points, across all particles.
Figure 9 provides a visual illustration of the pointwise VI contributions (VIC) when comparing particles 2 and 3 from Example 2. In this side-by-side plot, each data point is colored by its cluster assignment in the two particles, with shading indicating the magnitude of its contribution to the overall VI distance: solid points represent higher VIC values, while more transparent points indicate lower contributions. This visualization helps to pinpoint the specific data points (often located at the boundaries of clusters or those reassigned between clusters) that are most responsible for the differences between the two clustering solutions.
Figure 9: Side-by-side visualization of particles 2 and 3 from Example 2, with points colored by cluster assignment and shaded by their VI contribution (VIC). More solid points indicate higher VIC, highlighting areas where cluster assignments differ most between the particles; greater transparency corresponds to lower VIC and greater agreement.
VI Contribution by Group (VICG)
Recall from Section 3.2 that the VI contribution by group (VICG) aggregates the pointwise VI contributions for all data points within each cluster of the meet between two partitions. This provides a higher-level summary of how much each cluster in the meet contributes to the overall VI between two partitions. While individual pointwise VIC values can highlight subtle individual differences in clustering, the VICG effectively reveals more global structural changes, such as when a large cluster is split or merged differently across partitions.
Figure 10 visualizes the VICG for the comparison between particles 2 and 3 in Example 2. Unlike the pointwise VI contribution, which tends to emphasize small, local differences, the VICG highlights broader, high-level discrepancies between partitions. In this example, the meet clusters with the highest VICG, such as the top left component, correspond to large groups that undergo major changes—here, being merged with different components across the two particles. This groupwise perspective makes it easier to identify substantial structural differences in clustering solutions.
Figure 10: Visualization of VI contribution by group (VICG) between particles 2 and 3 in Example 2. The top left panel displays the cluster assignments defined by the meet of the two partitions, which in this case coincides with the meet of the three particles. The top right panel shows the VICG values for each meet cluster, quantifying the aggregate contribution of each group to the overall VI distance between the two partitions. The bottom panel provides a side-by-side comparison of the two particles, colored by cluster assignment and with transparency reflecting the VIC for each data point. Together, these panels highlight which subsets of data account for the largest discrepancies between the two clustering solutions.
Appendix D Experiments: additional material
In Section 4, we presented experiments with synthetic data under two scenarios where the posterior distribution displays high uncertainty: (1) when the model is correctly specified but the mixture components are not well-separated, and (2) when the model is misspecified. Here, we provide additional details on the data-generating processes for both scenarios.
For scenario (1), we investigated the impact of cluster separation on posterior uncertainty using datasets where the model is correctly specified but the mixture components have varying degrees of overlap. Each dataset consists of $n=600$ observations in two dimensions, drawn from a mixture of four equally weighted Gaussian components. The component means are symmetrically positioned at $(m,m)$, $(m,-m)$, $(-m,-m)$, and $(-m,m)$, where the separation parameter $m$ (with identity covariance matrix) controls the distance between clusters. The value of $m$ was varied across a dense grid to explore a wide range of overlap, specifically $m\in\{0.1,~{}0.3,~{}0.5,~{}0.7,~{}0.9,~{}1.1,~{}1.3,~{}1.5,~{}1.7,~{}1.9,~{}2.0,~{}2.2,~{}2.4,~{}2.6,$ 2.8,3.0}2.8,~{}3.0\}2.8 , 3.0 }. For each value of $m$, ten independent datasets were generated with different random seeds, with points assigned to each component with equal probability. A graphical representation is provided in Figure 11.
For scenario (2), we evaluated the robustness of WASABI under model misspecification, designing two types of synthetic experiments: (a) data drawn from a Gaussian distribution truncated to the unit square, and (b) data drawn from a two-dimensional skewed-$t$ distribution. In both cases, the aim is to investigate scenarios where a Gaussian mixture model is misspecified for the underlying data and to observe its effect on the posterior distribution and clustering summaries.
For each setting, datasets are generated as follows: for a given parameter configuration, $N=200$ data points in $p=2$ dimensions are simulated, and the process is repeated for five independent replicates. Each dataset is then analyzed using a (diagonal location-scale) Dirichlet process mixture of Gaussians fit via the BNPmix package (Corradin et al., 2021). Across both types, we explore a grid of parameter values to modulate the severity of misspecification.
[…]
For setting (b), each data point is drawn from a two-dimensional skewed-$t$ distribution, using the rskt function from the skewt package. Parameters are varied over a grid of degrees of freedom $df\in\{30,10,7.5,5,3\}$ and skewness parameters $\gamma\in\{1,1.5,2,3,5,10\}$, covering cases from nearly Gaussian to highly skewed and heavy-tailed. As before, datasets consist of $N$ points in two dimensions, and the Dirichlet process mixture model is fit using the same approach as in setting (a).
For each dataset generated under any setting, posterior draws are used to compute the posterior similarity matrix and clustering summaries, including minVI and the WASABI approximation.
Figure 11: Illustration for the data generating process for scenario (1) in the synthetic experiments. The data is generated from a four-components mixture Gaussian distribution, with each component centered around the points $(x_{i},y_{i})$ for $i=1,\ldots,4$, with $|x_{i}|=|y_{i}|=m$. The covariance matrix for each component is diagonal with marginal variances equal to 1. Several datasets are displayed for different values of the center location parameter $m$.
D.1 Posterior Compression
WASABI is useful for improved understanding of the posterior but could also be relevant for posterior compression in order to reduce both storage and the computational cost of predictive inference. For storage, saving only the WASABI particles and weights as opposed to all MCMC samples, reduces storage costs from $O(nT)$ to $O(nL)$, with $L<<T$. Predictive inference involves computing the predictive quantity of interest for every MCMC sample and all new points; for example, density estimation for multivariate Gaussian mixtures is $O(Tn^{*}p^{2}K_{\text{max}})$, where $n^{*}$ is the number of new points where the density is to be evaluated (often a large grid) and $K_{\text{max}}$ is the maximum number of clusters. If we are interested in conditional density estimation as in the HPV uptake study of Section 5, this becomes even more expensive as we must evaluate the density on grid of $y$ values for every covariate value considered.
!t
(a) MCMC density estimates
(b) WASABI density estimates
Figure 12: Density estimation (black) for Example 1 based on the MCMC samples and WASABI posterior, with the true data-generating density in red. For MCMC, the densities for a subset of MCMC draws are shown in gray, and for WASABI, the density for each particle is shown in gray with opacity determined by the particle’s weight.
Approximation method $L_{1}$ distance Wall-clock time (sec)
MCMC 0.114 1601.9
WASABI 0.139 1.7
Table 1: Density estimation for Example 1 based on an MCMC and WASABI approximation, comparing the tradeoff between low error (measured by the approximate $L_{1}$ distance between the true and estimated density) and run time.
For illustration, consider the task of density estimation for Example 1. While the MCMC-based estimates are slightly better than the WASABI-based ones (Figure 12), WASABI significantly reduces the cost by a factor of $1,000$ (Table 1).